OpenStudy (loser66):

What is difference between them? \(\sum_{n=0}^\infty \dfrac{n^n}{n!}x^n\) and \(\sum_{n=1}^\infty \dfrac{n^n}{n!}x^n\) I have to find the radius of the first one. I know the radius of the second one is 1/e but wonder how they are different. Appreciate any tips

OpenStudy (xapproachesinfinity):

well the first start at 0 and the second start at 1 so there is an extra term in the first

OpenStudy (loser66):

Does it affect the radius of convergence?

OpenStudy (xapproachesinfinity):

the second just starts at x

OpenStudy (xapproachesinfinity):

hmm i think not! you do the ratio test for both and you end up with the same radius of convergence

OpenStudy (xapproachesinfinity):

well in convergence i think they have the same behavior

OpenStudy (loser66):

kid, I have a very convenient formula to get it without apply ratio or root test. It is Hadamard theorem, but it applies to n=1, not to n=0, that is why made this question

OpenStudy (loser66):

I find lim sup of \(a_n\)

OpenStudy (xapproachesinfinity):

hmm i see

OpenStudy (xapproachesinfinity):

well perhaps it applicable because of theorems conditions but i think with other method the convergence should be the same, i think

OpenStudy (xapproachesinfinity):

mind showing how you do the entire convergence test with sup

OpenStudy (loser66):

One more question: If I use ratio test, what if I use \(\dfrac{a_n}{a_{n+1}}\)??

OpenStudy (loser66):

Ok

OpenStudy (loser66):

By Hadamard theorem, the radius of a convergence series is \(R =\dfrac{1}{lim_{n\rightarrow \infty}sup\sqrt[n]{|a_n|}}\)

OpenStudy (xapproachesinfinity):

hmm i don't that's allowable since we are going by \[0<a_n<b_n\] bn being a_n+1 here ratio test was a follow of comparison test

OpenStudy (xapproachesinfinity):

i might be wrong :)

OpenStudy (sdfgsdfgs):

sorry perhaps im dumb but i thought 0^0 is indetermine...so n cannot be zero if that is true?

OpenStudy (xapproachesinfinity):

hmm interesting theorem somehow it related to root test! can't tell further than that hehe since i didn't do much about the sups

OpenStudy (loser66):

\[ \sqrt[n]{|a_n|}=\dfrac{n}{\sqrt[n]{n!}}=\dfrac{n}{n/e}=e\]

OpenStudy (loser66):

hence R = 1/e

OpenStudy (loser66):

@sdfgsdfgs That is why I want to make it clear. :)

OpenStudy (loser66):

@xapproachesinfinity need the proof of \(\sqrt[n]{n!}= n/e\)??

OpenStudy (xapproachesinfinity):

hmm interesting but have some vague feeling to how you get n/e lol

OpenStudy (loser66):

hahahaha

OpenStudy (sdfgsdfgs):

@Loser66 if the first term at n=0 is indetermine, the whole summation will be indetermine as well...right? so there will be no convergence.

OpenStudy (xapproachesinfinity):

you read my mind lol had a suspicious feeling lol

OpenStudy (loser66):

@sdfgsdfgs Let discuss later, let me give @xapproachesinfinity the proof , ok?

OpenStudy (sdfgsdfgs):

@Loser66 sure :) grabbing popcorns (n my lunch! brb)

OpenStudy (xapproachesinfinity):

gotta go eat lunch, write the proof i will review it

OpenStudy (loser66):

ok, enjoy the meal, I will jot it down

OpenStudy (loser66):

\[ln(\sqrt[n]{n!})=\dfrac{1}{n}ln (n!)=\dfrac{1}{2}\sum_{k=1}^\infty ln(k)\] \[ln x\leq 2\in [1,2]\\lnx\leq 3\in [2,3]\\.....\\\int_1^n ln xdx\leq \sum_{k=1}^\infty ln(k)\]

OpenStudy (loser66):

\[\int_1^n lnx dx= xlnx -x|_1^n=nlnn-n+1\] \[ln (n!)\geq nln(n) -n+1\] on [1,2], \(ln x\geq ln1\) on [2,3], \(ln x\geq ln2\) .................................... \[\sum_{k=1}^n ln(x) =\sum_{k=1}^{n-1}ln k+ln n \leq \int_1^n lnx dx + lnx = nln n-n+1+ln n\]

OpenStudy (loser66):

\[nlnn-n+1\leq ln (n!)\leq nln n-n+1+ln n\\ln n-1+(1/n)\leq ln\sqrt[n]{n!}=(1/n)ln(n!)\leq ln n-1+(1/n)+(lnn/n)\] \[(1/n)\leq ln\sqrt[n]{n!}-ln(n+1)\leq \dfrac{1+lnn}{n}\]

OpenStudy (loser66):

as n--> infinitive, the far left and far right go to 0, by Squeeze theorem, the middle one goes to 0 also

OpenStudy (loser66):

Hence \[e^{ln\sqrt[n]{n!}-ln(n+1)} = middle~ term\]

OpenStudy (loser66):

\[e^{1/n}\leq\dfrac{{\sqrt[n]{n!}}}{n/e}\leq e^{1+lnn/n}\] hene, as n--> infinitive, the middle term goes to 1 by squeeze theorem, that gives \(\sqrt[n]{n!}=n/e\)

OpenStudy (loser66):

@sdfgsdfgs if n =0, then it is the \(\sum_{n=0}^\infty0^0\)

OpenStudy (loser66):

I do not know how to argue this case. Ha!!

OpenStudy (sdfgsdfgs):

@Loser66 nice proof above! but if n starts at 0, the first term kills the summation even thou the rest of terms converge. No way around it.

OpenStudy (loser66):

Ok, let me ask my prof on next Monday. If you are interesting still, remind me to give you the argument of my prof. :)

OpenStudy (sdfgsdfgs):

@Loser66 the summation beginning with n=0 is indetermine right? :)